Good morning is it possible to calculate when a dog stops accelerating with the
Dogs weight in kg
The distance traveled
And the time taken to travel
1
Expert's answer
2019-06-18T08:11:36-0400

Hello!

Yes, it is possible. The following expressions may help you. But first, we need to be sure what exactly we want to calculate about the dog.

Consider a dog with mass "m" which stops, i.e. decelerates. Assume that it had initial velocity "v_i" (fast) and final velocity "v_f" (slow, or the final speed is zero when the dog fully stops). Also, assume that the force of friction acting on the dog is "f". The acceleration (deceleration) will be hence


"a=\frac{f}{m}."

The distance traveled before the stop will be


"d=\frac{v_f^2-v_i^2}{2a}."

The time required for the dog to stop is


"t=\frac{v_f-v_i}{a}."